solve the following system of linear equations:
3x-2y+4z=1
x+y+2Z=3
2x-3y+6z=8

Answer by longjonsilver(2297)

well. This all depends what you know :-)

I would suggest Cramer's Rule, using determinants:

D, determinant of the main matrix = 20
Dx, determinant of main matrix with x-values substituted = -52
Dy, determinant of main matrix with y-values substituted = 12
Dz, determinant of main matrix with z-values substituted = 50

so x=Dx/D = -2.6

so y=Dy/D = 0.6

so z=Dz/D = 2.5

Put these values into the 3 equations and you get the correct answers.

Now, how do you calcualte the determinants? Ask me if this is your course. If you have no idea what i am gibbering on about, then ask again to have another solution method.

jon.